OpenStudy (anonymous):

plz help ( ; use benchmarks to estimate the sum 3 over 5 + 3 over 17 = ? a) 2 b) 1 over 2 c) 1 d) 0

jimthompson5910 (jim_thompson5910):

so this is \[\large \frac{3}{5} + \frac{3}{17}\] right?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

well, if we're just estimating, then 3/5 = 0.6 is close to 1 and 3/17 = 0.176 is close to 0 so (3/5) + (3/17) = 1 + 0 = 1 (based on the estimations above)

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

np